HUBBLE SPACE TELESCOPE DAILY REPORT #1051 PERIOD COVERED: 0700 EST Daily Status Report as of 012/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Six images were received and displayed in the OSS. Two of the images were reported to be underexposed, and the other four images were okay. 1.2 Completed FOS/BL/RD 4709 (SMOV FOS/FGS Coarse Alignment) The Faint Object Spectrograph (Blue and Red Sides) was used to make two sets of spectral observations of cluster NGC-199 to help determine the location of the 4.3 arcsecond aperture. The observations were received and displayed. Both images were reported to be normal. 1.3 Completed WF/PC-2 4616 (WF/PC-2 Fine Alignment and Focus) The WF/PC-2 was used to make observations of bright globular cluster Omega Centauri (Visual Magnitude 2-3, and Color Index ~0) both in the visual and ultraviolet, and using each of the CCDs at different focus settings. A total of 15 images was received and displayed. One image was reported to show the effects from a recentering event. The OSS reported that the effect of the secondary mirror movement was noticeable. 1.4 Completed FOC/96 4787 (SMOV FOC Throughput and Sensitivity) The Faint Object Camera (f/69) and COSTAR were used to make observations of the planetary nebula NGC-6543 to help measure the system's throughput. This also verifies that the COSTAR arm correctly blocks all of the direct light from reaching the FOC detector. As was expected, the single image showed no trace of the planetary nebula. 1.5 Completed WF/PC-2 5190 (Observations of External Galaxies) The WF/PC-2 was used to make images of a distant cluster of galaxies with a background quasi-stellar object. The OSS monitored the observations, and assessed the data quality. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled :9 Successful :9 2.3 Significant HSTARs: None 2.4 Operations Notes: The Secondary Mirror was successfully moved last night as part of the WF/PC-2 fine alignment (Prop. 4616). There were no reported problems in any onboard subsystem. All SMOV activities continue to be performed nominally. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibrations) 3.2 Complete Two Sets of FOS/RD/BL 4746 (COSTAR/FOS Coarse Alignment) 3.3 Complete WF/PC -2 5248 (WF/PC-2 Image Quality and Stability Check) 3.4 Complete Two Sets of WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn ON) 3.5 Complete FOS/BL 5133 (Mass Exchange Binaries (FOS 34) SS 2883: Cycle 4) 3.6 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-